An unknown silver/gray (92.9-g) piece of metal at 178.0°C is quickly placed to 75.0 mL of water initially at 24.0°C. After a few minutes, both the metal and the water have reached 29.7°C. Calculate the specific heat and the identity of the metal


1
Expert's answer
2021-11-08T08:11:37-0500

"TQ(H2O)=c(H2O)\u00d7\u03c1(H2O)\u00d7V(H2O)\u00d7\u0394T"

Q(H2O) = 1795.5 J

"c(Me) = \\frac{Q(H2O)}{m(Me) \\times \\Delta T1}\n\u200b"

"c(Me) = \frac{1795.5}{0.0929 \times (178-29.7)} = 130.33 \frac{J}{kg \times ^oC}"


.